# Quadratic exchange equations for Coxeter matroids

### v1

@@ 3.4.1 $E_{6}$ minuscule case
where the antipodes of $\lozenge(a_{i})$ are the sets $\left\{b_{j},c_{ij}\,\colon\,j\neq i\right\}$, the antipodes of $\lozenge(b_{i})$ are the sets $\left\{a_{j},c_{ij}\,\colon\,j\neq i\right\}$, and the antipodes of $\lozenge(c_{ij})$ are the sets $\{a_{i},b_{j}\},\{a_{j},b_{i}\}$, and $\{c_{k\ell},c_{k^{\prime}\ell^{\prime}}\}$ for all $\{k,k^{\prime},\ell,\ell^{\prime}\}=[6]\setminus\{ij\}$.
@@ 4.1.3 Matroids (type A)
Let $\mathbb{G}=\mathrm{SL}_{n}$ and $\mathbb{P}$ be the $k$-th maximal parabolic subgroup, so that $\mathbb{G}/\mathbb{P}$ is the Grassmannian $\mathrm{Gr}(k,n)$. The quadratic Grassmann–Plücker relations ${\mathcal{E}}^{(A)}_{n,k}$ are a basis for the equations of the embedding $\mathbb{G}/\mathbb{P}\subset\operatorname{proj}(\bigwedge^{k}{\mathbb{C}}^{n})$, and tropicalise to the type $A$ strong exchange equations ${\mathcal{F}}^{(A)}_{n,k}$.
@@ 4.2.1 Type $B$ embedding equations
$F^{(B)}_{I,J}$ $:=\sum_{i\in I\mathbin{\triangle}J}(-1)^{\ell\left(\sigma_{i,I}\right)+\ell\left(\sigma_{i,J}\right)}x_{I\mathbin{\triangle}i}x_{J\mathbin{\triangle}i}$ $\text{ if }|I\mathbin{\triangle}J|\equiv 0\mod 2$ (32)
$F^{(B)}_{I,J}$ $:=-x_{I}x_{J}+\sum_{i\in I\mathbin{\triangle}J}(-1)^{\ell\left(\sigma_{i,I}\right)+\ell\left(\sigma_{i,J}\right)}x_{I\mathbin{\triangle}i}x_{J\mathbin{\triangle}i}$ $\text{ if }|I\mathbin{\triangle}J|\equiv 1\mod 2$ (33)
@@
The claim in the theorem about tropicalisation is clear by inspection when one compares (3233) to (1415) above. We delay the proof of the rest until the end of this section.
For the proof, it will be convenient to rephrase and reindex these equations $F_{I,J}^{(B)}$. For disjoint $N,L\subseteq[n]$, define a face $Q_{N,L}=\left\{K\subset[n]\,\colon\,N\subset K,L\cap K=\varnothing\right\}$ of dimension n|N||L|n-|N|-|L| as a face of the $n$-cube $Q_{n}$. The full $n$-cube $Q_{n}$ is given by $Q_{\varnothing,\varnothing}$. Given a set $I$, write $I^{c}=[n]\setminus I$ for the complement. Any pair $(I,J)$ of subsets of $[n]$ is an antipode in a unique cube, namely $Q_{N,L}$, where $N=I\cap J$ and $L=(I\cup J)^{c}$ (this may be a 0-cube). So we will call such a pair of subsets an ordered antipode. Observe that we can write $J=(I^{c}\setminus L)\cup N$. As such, for any $K\in Q_{N,L}$ we let $\overline{K}_{N,L}:=(K^{c}\setminus L)\cup N$ denote the set such that $(K,\overline{K}_{N,L})$ is the unique antipode in $Q_{N,L}$ containing $K$. If $N$ and $L$ are clear from context we will drop them and simplify this to $\overline{K}$. Letting $S=(N\cup L)^{c}$, we have $\overline{K}_{N,L}=K\mathbin{\triangle}S$. With this notation fixed, we denote
We prove the following theorem in Appendix A.1. It is not a new result but we could not find a reference that produces these explicit quadratic polynomials. Recall the notation $\ell\left(-\right)$ for length of ordered multisets from Section 2.2. We let ${\mathcal{M}}_{N,L}$ denote the set
where m=n|N||L|m=n-|N|-|L|.
Theorem 4.9.
The set of equations
where
forms a basis for the Lichtenstein embedding equations.
Definition 4.10.
Given an ordered antipode $(K,\overline{K})$, we define
Recall that $N$ and $L$ are determined by $(K,\overline{K})$.
Observe that $K,\overline{K}]$ is independent of $M$. As such, with this notation the Lichtenstein embedding equations are simplified to
for all $M$ with $2|M|\in{\mathcal{M}}_{N,L}$. The proof of Theorem 4.9 shows that $p^{M}_{N,L}$ is symmetric under exchanging $K$ and $\overline{K}$, but in general $K,\overline{K}]$ differs from $\overline{K},K]$, and the two can even have different parities.
To prove Theorem 4.8 it suffices to show that the span of the quadratic embedding equations $\{\mathrm{ex}_{(K,K\mathbin{\triangle}S)}\}$ is equal to the span of the Lichtenstein embedding equations $\{p^{M}_{N,L}\}$. The rest of this subsection is dedicated to proving the equality of these subspaces of quadrics. Corollary 4.15 is the forward inclusion and Lemma 4.16 is the reverse.
To compare the quadratic embedding equations with the Lichtenstein embedding equations, we must rewrite them with coefficients in terms of $K,\overline{K}]$. This is the content of Corollary 4.12. We first give a technical lemma.
Lemma 4.11.
Proof.
We use Definition 4.10 to calculate the difference $K,\overline{K}]-[K\mathbin{\triangle}i,\overline{K}\mathbin{\triangle}i]\mod 2$ term by term. We first observe that
Repeating this argument for the other permutations in Definition 4.10, we get
Corollary 4.12.
Up to a global sign, we have
Proof.
If we multiply $\mathrm{ex}_{(K,K\mathbin{\triangle}S)}$ from Definition 4.7 by the sign $(-1)^{n+[K,K\mathbin{\triangle}S]}$, the claim immediately follows from Lemma 4.11. ∎
We next introduce the variables $y_{(K,\overline{K})}=(-1)^{[K,\overline{K}]}x_{K}x_{\overline{K}}$. We emphasise that as $(-1)^{[K,\overline{K}]}$ may differ from $(-1)^{[\overline{K},K]}$, the variables $y_{(K,\overline{K})}$ and $y_{(\overline{K},K)}$ may differ. By Corollary 4.12, in these variables, the quadratic embedding equations become
and the Lichtenstein embedding equations become
This change of variables helps us prove the following.
Lemma 4.13.
Proof.
We next define our intermediary polynomials $\chi_{b}$, as follows:
$\chi_{b}=\sum_{|M|=b}p^{M}_{\varnothing,S^{c}}(y)\,.$
We prove (40) by instead considering a more general function
$S(m,a,c)\coloneqq\sum_{b\equiv c\bmod 2}\frac{1}{2}(c-b)\chi_{b}(m,a)$
where $c\in{\mathbb{Z}}$ and
$\chi_{b}(m,a):=\sum_{k}(-1)^{k}\binom{a}{k}\binom{m-a}{b-k}\,$.
$S(m,0,c)=\frac{1}{4}(S^{+}(m,c)+(-1)^{c}S^{-}(m,c)).$
When $a>0$ we have
$=\chi_{b}(m-1,a-1)-\chi_{b-1}(m-1,a-1)$
We finish by doing this check. When $c$ is even,
$S(3,a,c)$ $=\frac{c}{2}\chi_{0}(3,a)+\frac{c-2}{2}\chi_{2}(3,a)$
$=\frac{c}{2}+\frac{c-2}{2}(2a^{2}-6a+3).$
The evaluations of $a$ appearing in (41) are
$S(3,0,c)$ $=2c-3=2^{3-2}(c-\textstyle\frac{3}{2}),$
$S(3,1,c)$ $=1=2^{3-3}.$
In the same way, when $c$ is odd,
$S(3,a,c)$ $=\frac{c-1}{2}\chi_{0}(3,a)+\frac{c-3}{2}\chi_{2}(3,a)$
$=\frac{c-1}{2}(-2a+3)+\frac{c-3}{2}(-\frac{4}{3}a^{3}+6a^{2}-\frac{20}{3}a+1),$
and the evaluations in (41) are again
$S(3,0,c)$ $=2c-3,$
$S(3,1,c)$ $=1$
as desired. ∎
Having shown that every $\operatorname{ex}_{(\varnothing,S)}$ is in the span of the Lichtenstein embedding equations, the next step is to generalise to all quadratic embedding equations. Every antipode of every subcube of the $n$-cube is in the orbit of an antipode $\{\varnothing,S\}$ under reflections in the coordinate hyperplanes, so our strategy is essentially to apply these reflections. But we cannot literally act by symmetries of the cube; the symmetries we have are those of the spin representation $\mathrm{Spin}(2n+1)$, worked with in Appendix A.1. We must therefore briefly introduce a little of that representation theory here: for full details see the appendix.
Let $V={\mathbb{C}}^{2n+1}$ be the vector representation of $\mathfrak{so}_{2n+1}$ with non-degenerate bilinear form $B$. There exists a pair of maximal isotropic subspaces $F$, $F^{\vee}$ with respect to $B$ and a unit length vector $w\in V$ orthogonal to $F$ and $F^{\vee}$ such that $V=F\oplus F^{\vee}\oplus\mathrm{Span}\{{w}\}$. We have $B(w,w)=1$, and moreover we fix bases $\{v_{1},\dots,v_{n}\}$ for $F$ and $\{x_{1},\dots,x_{n}\}$ for $F^{\vee}$ such that $B(v_{i},x_{j})=\delta_{ij}$. With this, we model $\mathcal{S}^{\vee}={\textstyle{\bigwedge}}^{\bullet}F^{\vee}$ with basis $\left\{x_{I}:=x_{i_{1}}\wedge\cdots\wedge x_{i_{k}}\,\colon\,I=\{i_{1},\dots,i_{k}\}\subseteq[n]\right\}$.
$\frac{1}{\sqrt{2}}(x_{i}+v_{i})w\in\mathrm{Spin}(2n+1)\,.$
Thus in both cases, in the notation of Section 2.2, we have
Since $g_{i}$ is a group element, we have
$g_{i}(x_{A}x_{B})=g_{i}(x_{A})g_{i}(x_{B})\,.$
Lemma 4.14.
Proof.
For any three sets $A,B,C$, we have the associativity relation
Thus, for any $A\subseteq[n]$, we have
Corollary 4.15.
Proof.
This follows from Lemma 4.14 because the span of this set of $p^{M}_{N,L}$ is closed under the action of the $g_{h}$. ∎
Lemma 4.16.
Proof.
$p^{M}_{N,L}[x_{I}x_{\overline{I}}]=(-1)^{|I^{c}\cap M|+[I,\overline{I}]}.$
Note that
Equivalently,
Hence, we find
We also recall from Lemma 4.11 that
$=(-1)^{|I^{c}\cap M|+[I,\overline{I}]}\,.$
Proof of Theorem 4.8.
Corollary 4.15 and Lemma 4.16 prove that the span of the quadratic embedding equations $\{\mathrm{ex}_{(K,K\mathbin{\triangle}S)}\}$ is equal to the span of the Lichtenstein embedding equations $\{p^{M}_{N,L}\}$. As $\{p^{M}_{N,L}\}$ cut out ${\mathbb{G}}/{\mathbb{P}}\subset\operatorname{proj}(\mathcal{S})$ by Theorem 4.9, our theorem is proved. ∎
@@
Unlike the equations in Theorem 4.9, we emphasise that the type $B$ quadratic embedding equations do not form a basis for the Lichtenstein embedding equations, as the following example demonstrates. Consider the $n=5$ case: this corresponds to the minuscule variety $\mathrm{Spin}(11)/{\mathbb{P}}_{\alpha_{5}}\subseteq\operatorname{proj}(\mathcal{S})$ where $\dim(\mathcal{S})=2^{5}$. The analysis in Proposition A.9 demonstrates that $S^{2}(\mathcal{S})$ decomposes into ${\textstyle{\bigwedge}}^{2}{\mathbb{C}}^{11}\oplus{\textstyle{\bigwedge}}^{6}{\mathbb{C}}^{11}\oplus{\textstyle{\bigwedge}}^{10}{\mathbb{C}}^{11}$, where ${\textstyle{\bigwedge}}^{6}{\mathbb{C}}^{11}$ is the component with highest weight $2\lambda=(1,\dots,1)$. It follows from Lemma 4.3 that the space of Lichtenstein embedding equations is of dimension $\binom{11}{2}+\binom{11}{10}=66$. Moreover, the space of Lichtenstein embedding equations of weight zero is six-dimensional, corresponding to the direct sum of the Cartan subalgebra (dimension $5$) and the $1$-dimensional span of the weight zero vector in $({\mathbb{C}}^{11})^{\vee}\cong{\textstyle{\bigwedge}}^{10}{\mathbb{C}}^{11}$. This can be equivalently computed by counting subsets $M\subseteq[5]$ such that
@@
On the other hand, the quadratic embedding equations of weight zero are precisely those $F_{I,J}$ with $I\mathbin{\triangle}J=[5]$. The number of these is equal to the number of antipodes, which is $2^{4}=16$. This demonstrates that, although the quadratic embedding equations span the Lichtenstein embedding equations, they are far from a basis.
@@ 4.4.1 $E_{6}$ minuscule case(s)
To explicitly write these equations, set $X=\left\{x_{A}\,\colon\,A\in W(E_{6})^{S\setminus s_{6}}\right\}$ where we recall that $S=\{s_{1},\dots,s_{6}\}$ is the set of simple reflections for $W(E_{6})$. Recalling that each coset $A$ has an associated minimal coset representative $w_{A}$, we can equivalently think of $X$ as variables, or as a basis for $V_{\lambda}^{\vee}$, where $x_{A}:=x_{w_{A}(\lambda)}$ is the vector associated to weight $w_{A}(\lambda)$. Consider the equation
@@ 4.4.2 $E_{7}$ minuscule case
To explicitly write these equations, set $X=\left\{x_{A}\,\colon\,A\in W(E_{7})^{S\setminus s_{7}}\right\}$ where we recall that $S=\{s_{1},\dots,s_{7}\}$ are the set of simple reflections for $W(E_{7})$. Recalling that each coset $A$ has an associated minimal coset representative $w_{A}$, we can equivalently think of $X$ as variables, or as a basis for $V_{\lambda}^{\vee}$, where $x_{A}:=x_{w_{A}(\lambda)}$ is the vector associated to weight $w_{A}(\lambda)$. Consider the equation
@@ A.2 Explicit equations: Type $B_{n}$
We now have an explicit description of the decomposition of $S^{2}(\mathcal{S}^{\vee})$ into irreducible $\mathfrak{so}_{2n+1}$-modules via the four isomorphisms described in Lemma A.7, Lemma A.8, Corollary A.5 and Lemma A.2. This subsection is dedicated to the slightly tedious process of carefully pulling a basis for ${\textstyle{\bigwedge}}^{\bullet}V$ back through these three isomorphisms.
Recall the following notation from Section 4.2.1. We let $Q_{N,L}=\left\{K\subset[n]\,\colon\,N\subset K,\,L\cap K=\varnothing\right\}$ denote the set of sets located on a face of our cube $Q_{n}$. Recall that for $K\in Q_{N,L}$, we let $\overline{K}:=(K^{c}\setminus L)\cup N$ denote its antipode in $Q_{N,L}$. Additionally, as in (35) we let
${\mathcal{M}}_{N,L}=\left\{4k+2\left\lceil\frac{n}{2}\right\rceil-|N|-|L|-\varepsilon\,\colon\,k\in\mathbb{Z}\,,\,\varepsilon\in\{0,1\}\right\}\setminus\Big\{m-2,\,m-1,\,m,\,m+1,\,m+2\Big\}$
where m=n|N||L|m=n-|N|-|L| is the dimension of the cube $Q_{N,L}$.
See 4.9

### v2

@@ 3.4.1 $E_{6}$ minuscule case
where the antipodes of $\lozenge(a_{i})$ are the sets given by $\left\{b_{j},c_{ij}\,\colon\,j\neq i\right\}$, the antipodes of $\lozenge(b_{i})$ are the sets given by $\left\{a_{j},c_{ij}\,\colon\,j\neq i\right\}$, and the antipodes of $\lozenge(c_{ij})$ are the sets given by $\{a_{i},b_{j}\},\{a_{j},b_{i}\}$, and $\{c_{k\ell},c_{k^{\prime}\ell^{\prime}}\}$ for all $\{k,k^{\prime},\ell,\ell^{\prime}\}=[6]\setminus\{ij\}$.
@@ 4.1.3 Matroids (type A)
Let ${\mathbb{G}}=\mathrm{SL}_{n}$ and $\mathbb{P}$} be the $k$-th maximal parabolic subgroup, so that $\mathbb{G}$$\mathbb{P}$} is the Grassmannian $\mathrm{Gr}(k,n)$. The quadratic Grassmann–Plücker relations ${\mathcal{E}}^{(A)}_{n,k}$ are a basis for the equations of the embedding ${\mathbb{G}}/{\mathbb{P}}\subset\operatorname{proj}(\bigwedge^{k}{\mathbb{C}}^{n})$, and tropicalise to the type $A$ strong exchange equations ${\mathcal{F}}^{(A)}_{n,k}$.
@@ 4.2.1 Type $B$ embedding equations
$F^{(B)}_{I,J}$ $:=\sum_{i\in I\mathbin{\triangle}J}(-1)^{\ell\left(\sigma_{i,I}\right)+\ell\left(\sigma_{i,J}\right)}x_{I\mathbin{\triangle}i}x_{J\mathbin{\triangle}i}$ $\text{ if }|I\mathbin{\triangle}J|\equiv 0\mod 2\phantom{.}$ (32)
$F^{(B)}_{I,J}$ $:=-x_{I}x_{J}+\sum_{i\in I\mathbin{\triangle}J}(-1)^{\ell\left(\sigma_{i,I}\right)+\ell\left(\sigma_{i,J}\right)}x_{I\mathbin{\triangle}i}x_{J\mathbin{\triangle}i}$ $\text{ if }|I\mathbin{\triangle}J|\equiv 1\mod 2.$ (33)
@@
The claim in the theorem about tropicalisation is clear by inspection when one compares (3233) to (1415) above. The proof of the rest is the subject of Sections A.1 and A.2. The equations are not new, but we have provided a proof as we could not find a reference that produces these explicit quadratic polynomials, especially with their signs overt. See further Remark 4.12.
@@
We emphasise that the type $B$ quadratic embedding equations do not form a basis for the Lichtenstein embedding equations, as the following example demonstrates. Consider the $n=5$ case: this corresponds to the minuscule variety $\mathrm{Spin}(11)/{\mathbb{P}}_{\alpha_{5}}\subseteq\operatorname{proj}(\mathcal{S})$ where $\dim(\mathcal{S})=2^{5}$. The analysis in Proposition A.9 demonstrates that $S^{2}(\mathcal{S})$ decomposes into ${\textstyle{\bigwedge}}^{2}{\mathbb{C}}^{11}\oplus{\textstyle{\bigwedge}}^{6}{\mathbb{C}}^{11}\oplus{\textstyle{\bigwedge}}^{10}{\mathbb{C}}^{11}$, where ${\textstyle{\bigwedge}}^{6}{\mathbb{C}}^{11}$ is the component with highest weight $2\lambda=(1,\dots,1)$. It follows from Lemma 4.3 that the space of Lichtenstein embedding equations is of dimension $\binom{11}{2}+\binom{11}{10}=66$. Moreover, the space of Lichtenstein embedding equations of weight zero is six-dimensional, corresponding to the direct sum of the Cartan subalgebra (dimension $5$) and the $1$-dimensional span of the weight zero vector in $({\mathbb{C}}^{11})^{\vee}\cong{\textstyle{\bigwedge}}^{10}{\mathbb{C}}^{11}$. This can be equivalently computed via the analysis in Section A.2.1 by counting subsets $M\subseteq[5]$ such that
@@
On the other hand, the quadratic embedding equations of weight zero are precisely those $F_{I,J}^{(B)}$ with $I\mathbin{\triangle}J=[5]$. The number of these is equal to the number of antipodes, which is $2^{4}=16$. This demonstrates that, although the quadratic embedding equations span the Lichtenstein embedding equations, they are far from a basis.
Note that in Theorem A.10 we give equations that do form a basis for the Lichtenstein embedding equations.
@@ 4.2.2 Type $D$ embedding equations
Remark 4.12.
The Type $B_{n}$ minuscule variety, the orthogonal Grassmannian $\mathrm{OGr}(n,2n+1)$, is diffeomorphic to either of the type $D_{n+1}$ minuscule varieties, the orthogonal Grassmannians $\mathrm{OGr}^{+}(n+1,2n+2)\cong\mathrm{OGr}^{-}(n+1,2n+2)$. In fact either of the half spin representations of $\mathrm{Spin}(2n+2)$, restricted to $\mathrm{Spin}(2n+1)$, is isomorphic to the single spin representation. Thus one can include the spin representation of $\mathrm{Spin}(2n+1)$ in either of the half spin representations of $\mathrm{Spin}(2n+2)$ (Lemma A.20). The isomorphism sends the type $B_{n}$ variable $x_{I}$ to the type $D_{n+1}$ variables $x_{I}$ or $\sqrt{-1}x_{I\cup\{n+1\}}$ depending on which variable has a subscript of the correct parity. Hence, this induces an isomorphism between the Lichtenstein embedding equations of $\mathrm{OGr}(n,2n+1)$ and the Lichtenstein embedding equations of $\mathrm{OGr}^{\pm}(n+1,2n+2)$, and likewise between the two sets of the quadratic embedding equations. Note that the map from Remark 3.10 is the combinatorial analogue of this isomorphism.
In fact the equations ${\mathcal{E}}_{n,+}^{(D)}$ are the Wick relations [conciniprocesi82]. We failed to find a reference that proves the Wick relations cut out the orthogonal Grassmannian $\mathrm{OGr}^{+}(n+1,2n+2)$ although this does seem to be folklore, mentioned in [sturmfelsvelasco2010] and [Rincon]. Manivel gives these equations up to signs in [manivel2009, 6], in which he states
The [signs] can be written down explicitly but seemingly not in a pleasant way. These relations are probably known, but we do not know any suitable reference.
This reflects our feeling precisely. Procesi states only that these equations can be deduced from an equation [procesi2007, 434] using two identities [procesi2007, (7.2.3)] and [procesi2007, (7.2.1)]. Seshadri [seshadri1978geometry] gives quadratic equations up to signs with different support. In [chirivi2013pfaffians, Corollary 16], Chirivì and Maffei give an alternative presentation using shuffle relations, but their equations have strictly larger support. It seems that in the literature the equations ${\mathcal{E}}_{n}^{(D)}$ are implicitly mentioned but not explicitly calculated. For completeness we have therefore included a full derivation.
@@ 4.4.1 $E_{6}$ minuscule case(s)
To explicitly write these equations, set $X=\left\{x_{A}\,\colon\,A\in W(E_{6})^{S\setminus s_{6}}\right\}$ and recall that $S=\{s_{1},\dots,s_{6}\}$ is the set of simple reflections for $W(E_{6})$. Recalling that each coset $A$ has an associated minimal coset representative $w_{A}$, we can equivalently think of $X$ as variables, or as a basis for $V_{\lambda}^{\vee}$, where $x_{A}:=x_{w_{A}(\lambda)}$ is the vector associated to weight $w_{A}(\lambda)$. Consider the equation
@@ 4.4.2 $E_{7}$ minuscule case
To explicitly write these equations, set $X=\left\{x_{A}\,\colon\,A\in W(E_{7})^{S\setminus s_{7}}\right\}$ and recall that $S=\{s_{1},\dots,s_{7}\}$ are the set of simple reflections for $W(E_{7})$. Recalling that each coset $A$ has an associated minimal coset representative $w_{A}$, we can equivalently think of $X$ as variables, or as a basis for $V_{\lambda}^{\vee}$, where $x_{A}:=x_{w_{A}(\lambda)}$ is the vector associated to weight $w_{A}(\lambda)$. Consider the equation
@@ A.2 Explicit equations: Type $B_{n}$
This section is divided into two parts. In the previous section we obtained an explicit description of the decomposition of $S^{2}(\mathcal{S}^{\vee})$ into irreducible $\mathfrak{so}_{2n+1}$-modules via the four isomorphisms described in Lemma A.7, Lemma A.8, Corollary A.5 and Lemma A.2. Section A.2.1 is dedicated to the slightly tedious process of carefully pulling a basis for ${\textstyle{\bigwedge}}^{\bullet}V$ back through the three isomorphisms, culminating in Theorem A.10 below. Then, in Section A.2.2, we prove Theorem 4.8 by confirming that its equations have the same span as those in Theorem A.10.
A.2.1 A first basis of equations
We prove Theorem A.10 in this section. Recall the notation $\ell\left(-\right)$ for length of ordered multisets from Section 2.2. We let ${\mathcal{M}}_{N,L}$ denote the set
where m=n|N||L|m=n-|N|-|L|.
For disjoint $N,L\subseteq[n]$, define the set $Q_{N,L}=\left\{K\subset[n]\,\colon\,N\subset K,L\cap K=\varnothing\right\}$ of dimension n|N||L|n-|N|-|L|. This is the set of subsets indexing a face of the $n$-cube $Q_{n}$; we’ll generally speak of it as a face or a (sub)cube itself. The full $n$-cube $Q_{n}$ is given by $Q_{\varnothing,\varnothing}$. Given a set $I$, write $I^{c}=[n]\setminus I$ for the complement. Any pair $(I,J)$ of subsets of $[n]$ is an antipode in a unique cube, namely $Q_{N,L}$, where $N=I\cap J$ and $L=(I\cup J)^{c}$ (this may be a 0-cube). So we will call such a pair of subsets an ordered antipode. Observe that we can write $J=(I^{c}\setminus L)\cup N$. As such, for any $K\in Q_{N,L}$ we let $\overline{K}_{N,L}:=(K^{c}\setminus L)\cup N$ denote the set such that $(K,\overline{K}_{N,L})$ is the unique antipode in $Q_{N,L}$ containing $K$. If $N$ and $L$ are clear from context we will drop them and simplify $\overline{K}_{N,L}$ to $\overline{K}$. Letting $S=(N\cup L)^{c}$, we have $\overline{K}_{N,L}=K\mathbin{\triangle}S$.
Theorem A.10.
The set of equations
where
forms a basis for the Lichtenstein embedding equations.